### v1

# High-performance superconducting two-qubit gate using a three-mode tunable coupler

Abstract
[…]
I Introduction
Development of superconducting qubit-based quantum computing devices has achieved a number of major milestones during the past few years [1, 2, 3, 4, 5]. The common theme is the usage of transmon qubits [6] with dispersive readout and capacitive coupling [7]. Nevertheless, the relatively large two-qubit gate error rates undermine both the practical application of such machines as well as the proof-of-principle advantage demonstration [8].
For implementation of two-qubit gates, there has been a variety of different approaches, each with its strengths and weaknesses. The simplest scheme involves resonant exchange of qubit populations triggered by a frequency shift of one of the qubits [9, 10]. For transmons, one can choose between an iSWAP-type gate [11, 12, 13] or a CPHASE-type gate [12, 13, 14]. The gate duration can be short and is determined by the qubit-qubit coupling constant. This scheme is prone to (i) significant qubit dephasing, because the qubits are operated outside of their sweet spots, which can be addressed with net-zero pulses [15], (ii) crossing of two-level systems (TLS) defects during gate operation [16], and (iii) large residual ZZ crosstalk when the qubits are detuned from each other. The need to traverse a large frequency range with the qubit and thus potentially cross strongly coupled TLS defects can be circumvented, for example, using a parametrically modulated signal [17]. For parametrically activated gates, however, the effective coupling strength is only a fraction of the underlying coupling between qubits. Moreover, the qubit is still detuned from it’s sweet spot, resulting in additional dephasing, which can be reduced by applying a modulation around the DC sweet spot [18].
Another approach is by using a tunable coupler element, either for the parametric modulation gate [19, 20] or direct resonance gate [21, 22]. For CPHASE-type gates one can also use an adiabatic flux pulse, which results in slower gates for the same coupling strength [23, 24, 25, 26, 27], but does not require precise resonance between the qubits. Otherwise, one must tune both the qubit frequency and the coupler. To avoid dephasing, one can use an AC signal to tune the effective qubit frequency for the duration of the gate [28]. By controlling both qubit frequencies and coupler flux with DC signals, one can implement the full continuous gate set that, up to single-qubit gates, includes all excitation-number-preserving unitaries [12]. For tunable couplers, one of the main desired features is cancelling out crosstalk. This can be done in a straightforward manner by using destructive interference between two coupling paths. These coupling paths can either be of Fano-type, with one coupling being resonant [29] or both coupling paths having resonances [30]. In either case, the residual interaction strongly depends on the relation between qubit and coupler frequencies.
In this work, we present a unitary cell suitable for a scalable quantum processor architecture and consisting of a two concentric transmons [31, 32, 17, 33, 34] and a tunable three-mode coupler for a controlled ZZ interaction. The coupler consists of a coplanar shorted by a SQUID in the middle and Josephson junctions (JJs) at the coplanar ends, being an upgrade of a promising tunable coupler [30, 35, 29]. Two JJs at the coplanar ends are implemented in order to suppress a parasitic ZZ interaction. Thus, by tuning the magnetic flux in the coupler SQUID, we are able to vary the interaction strength between qubits by two orders of magnitude. A native two-qubit CZ gate is implemented by leveraging a cross-Kerr interaction [30, 36, 35, 29, 37]. Our experimental CZ cross-entropy benchmarking (XEB) repeatedly gives the gate fidelity above 98%, being limited mostly by the qubits coherence time.
II Device description
The two-qubit device explored in this work is shown in Fig. 1. It consists of two tunable transmons connected by a tunable three-mode coupler with an individual flux-control line. We utilize here concentric weakly flux-tunable transmon qubits [38] for reduction of dielectric losses, containing a 3-JJs circuit [39] for reducing flux noise sensitivity at the expence of narrowing tunability range, shown in Fig. 1. Our 3-JJs qubit circuit consists of two JJs forming an asymmetric DC SQUID connected in series to a third JJ. Each qubit has a single flux-coupled control line for both dc and microwave manipulations. The physical pitch between qubits is 2 mm. Capacitive coupling between the transmon and the coupler is provided by arc-shaped electrodes [17, 40]. The three-mode coupler, that we present in the current work, has three nodes and consists of a coplanar waveguide segment, terminated with JJs, and shorted in the middle by a DC SQUID, see Fig. 1(b).
The Hamiltonian of the whole qubit-coupler-qubit system can be written as:
$\hat{H}_{\mathrm{sys}}=\sum_{i\in\{{Q_{\mathrm{A}},Q_{\mathrm{B}},c_{1},c_{2},c_{3}\}}}\hat{H}_{i}+\hat{H}_{\text{int}},$ (1)
where we consider our system as five subsystems associated with 5 nodes in Fig. 1(c). Two of nodes $Q_{\mathrm{A}}$ and $Q_{\mathrm{B}}$ correspond to the qubits, while the rest three $c_{1},c_{2},c_{3}$ refer to the coupler modes. The interaction term takes the following form:
$\hat{H}_{\text{int}}=\sum_{i\neq j}E_{Cij}\hat{n}_{i}\hat{n}_{j},$ (2)
where $\{i,j\}$ is the set of circuit nodes pairs and $E_{Cij}$ is the energy of the capacitive interaction. Exact expression of the Hamiltonian (1) is given in Appendix A.
Figure 1: The two-qubit cell. (a) An optical image of a sample. Each transmon (red) has a round island grounded through a circuit of three Josephson junctions: one junction in series with a SQUID (yellow). Each qubit has a separate control line both for tuning and excitation. The transmons are connected to each other through a tunable coupler (blue). The coupler shares an arc-shaped mutual capacitor with each qubit. The coupler is realised by a coplanar line with a SQUID in the middle. A separate line controls a magnetic flux through coupler’s SQUID providing tunability. In order to suppress parasitic ZZ interaction, an additional Josephson junction in each end of the coplanar is embedded. (b) Three key elements of the cell (from left to right): the three Josephson junction circuit of the transmon, the Josephson junction at one of the coplanar ends, the SQUID of the coupler. (c) A principal electrical circuit of the two-qubit cell. The two transmons are shown in red, the coupler is shown in blue. The more detailed electrical circuit can be found in Appendix A.
Figure 2: The measurement of the ZZ interaction. (a) The pulse sequence performed on our device: a Ramsey-type experiment on the qubit B while control qubit A is initialized in its ground or excited state, the coupler is tuned by a Hann-shaped flux pulse of the duration $t_{d}$. The argument $\varphi$ of the virtual $R_{Z}$ gate is determined by an offset frequency $f_{\text{offset}}$ and pulse duration $t_{d}$. (b, c) The resulted pattern of the qubit B relative excitation. The red line indicates the boundary between two different offset frequencies used in the experiments. (d) The ZZ interaction strength defined as the difference between the oscillation frequencies of the qubit B population in (b) and (c). The red dashed line shows the ZZ interaction strength $\zeta_{ZZ}$, calculated using the full-circuit model described in Appendix A with the experimental parameters obtained by two-tone spectroscopy measurements of each computational qubit and the coupler. The optimized parameters are given in Appendix A Table 3.
III Results
III.1 ZZ interaction strength
During the study of the ZZ interaction between qubits, both transmons are kept in the upper flux sweet spot, while the magnetic flux in the coupler loop is varied between zero flux and half flux quantum $\Phi_{\mathrm{ext}}\in\left[0,\Phi_{0}/2\right]$. In each subsystem, we take into account the first five energy levels. By numerical diagonalization of the system Hamiltonian (1) we analitically calculate a ZZ-interaction strength shown in Fig. 2(d) together with experimental data. Details on the numerical calculation can be found in Appendix A.
In our experiment we measure the strength of the ZZ-interaction between qubits as a function of an external flux through the coupler SQUID. We use a Ramsey-type pulse sequence, consisting of two $\pi/2$ pulses applied to the qubit B, and a Hann-shaped flux pulse of duration $t_{d}$, applied to the coupler, between them, as depicted in Fig. 2(a). We activate [Fig. 2(c)] or do not activate [Fig. 2(b)] the ZZ interaction with an $X$ operation on the control qubit A and measure the population of the target qubit B. For each voltage amplitude generating a flux through coupler SQUID, we find the frequency of population oscillations as a function of the delay $t_{d}$ for both cases when the qubit A remains at the ground state [Fig. 2(b)] and goes to the excited state [Fig. 2(c)]. We define the strength of the ZZ interaction $\zeta_{ZZ}$ as the difference of these frequencies. At this point we should mention, that in order to observe oscillations on scales of hundreds of nanoseconds even at low external flux, we apply virtual rotation $R_{Z}(\varphi)$ [41] about Z axis on the qubit B before the second $R_{X}(\pi/2)$ gate. The phase $\varphi$ depends on the duration of the coupler pulse $\varphi=2\pi f_{\text{offset}}t_{d}$, where $f_{\text{offset}}$ is a defined offset from the frequency of the qubit B. For convenience, we take two offset frequencies equal to 10 MHz for low and 50 MHz for high CZ-pulse amplitudes. The border of the offset frequency change is denoted with the red line in Fig. 2(b, c).
We provide the measured ZZ interaction strength in Fig. 2(d) as blue points. We extract system parameters from the two-tone spectroscopy of both qubits. Knowing the qubit frequencies $f_{\text{A}},f_{\text{B}}$ and anharmonicities $\alpha_{\text{A}},\alpha_{\text{B}}$ as functions of qubit external flux, one can get a critical current of two 3-JJs circuits of the qubits, marked with the red color in Fig. 1. From the two-tone spectroscopy of the coupler we obtain the coupler frequency $f_{\text{C}}$ at external magnetic flux $\Phi_{\mathrm{ext}}=0.5\Phi_{0}$. Using the full system Hamiltonian with obtained qubit JJs parameters and capacitances from the design, we fit the four JJs related to the coupler, colored blue in Fig. 1(c). We optimize the three critical currents of the coupler junctions in order to meet an experimental values of the ZZ interaction at the zero external magnetic flux $\zeta_{ZZ}(\Phi_{\text{ext}}=0)$ and the coupler frequency $f_{\text{C}}$. The designed and fitted parameters required for the system Hamiltonian (1) are presented in Tables 2 and 3 correspondingly. Also, we should mention that all junctions were produced in one fabrication stage, having the same critical current density. Thus, during the optimization, we vary only the JJs areas, assuming the same critical current density for all of them. Finally, the calculated the ZZ interaction strength as a function of the external magnetic flux $\Phi_{\text{ext}}$ plotted in Fig. 2(d).
The residual ZZ interaction at $\Phi_{\text{ext}}=0$, i.e. when the coupler is deactivated, was obtained to be 416 kHz, which is almost 300 kHz higher then expected. The three-mode tunable coupler calculated for the designed parameters gives the ZZ interaction strength $\zeta_{ZZ}$ below 150 kHz (see Fig. 7(b) in Appendix A). Such a difference in $\zeta_{\mathrm{ZZ}}$ is caused by fabrication imperfections, i.e. by the difference between the design and the fabricated device. However, the difference between expected and approximated parameters does not significantly influence the overall gate performance, as we show in further simulations and experiments.
Figure 3: The CZ gate simulation for the experimentally obtained device parameters. (a) The fidelity $F$ (red), the target population $P_{t}$ calculated as a mean projection of the computational states on themselves during the CZ gate (dashed blue) and the high-energy levels leakage error $L_{e}$ (dashed-dotted green) are shown as functions of the CZ gate duration. (b) The corresponding amplitude of the external flux in the coupler’s SQUID. (c) For the gate duration of 43.3 ns (star in panel (a)) the time evolution of the conditional two-qubit phase $\theta$ (red), the target population $P_{t}$ (dashed blue), and the leakage error $L_{e}$ (dashed-dotted green) are simulated. The fidelity estimated by Eq. (3) is $F=0.999(7)$. (d) The coupler flux pulse with Hann-shaped edges is used both in simulations and experiments.
III.2 CZ gate concept
Since changing the external magnetic flux in the coupling element, directly controls the ZZ interaction between qubits, the idea of the controlled-phase gate implementation looks quite natural. One can switch on the interaction adiabatically increasing the external flux and, after waiting for the required time, return the coupler to the initial state. The flux pulse is supposed to be short while remaining adiabatic in order to prevent undesired excitation transitions between the computational states and leakage out of the computational subspace. Therefore, in order to activate the interaction we use a smooth pulse $\Phi_{\text{ext}}(t)$ with a Hann function-shaped edges of 9 ns both in the simulation and the experiment, as it is shown in Fig. 3(d). Also, we define the gate duration as the length of the entire pulse, accounting both edges and a plateau-shaped part.
We described above in details the method of constructing the system model using two-tone spectroscopy data and measurement of the ZZ-interaction strength at the zero-flux point. Using the obtained values, we construct the full-circuit Hamiltonian and calculate the time dynamics of the system during the CZ gate implementation. In this numerical analysis, we take into account the first four energy levels of each qubit and each mode of the coupler in order to reduce a computational time. Thus, the Hilbert space dimension of the full system in the simulation is $4^{5}$. We denote the states of each computational qubit as $|g\rangle,|e\rangle,|f\rangle$ and use numerical notation for coupler modes.
We simulate the CZ gate performance for an amplitude of the coupler’s SQUID flux pulse in a range from $0.37\Phi_{0}$ to $0.5\Phi_{0}$. For each amplitude value we find a gate duration, making the state $|ee\rangle$ accumulated conditional phase $\theta=\pi$. In order to evaluate the performance of the resulting CZ operation, we use the standard expression for the two-qubit gate fidelity [42, 43]:
$F=\frac{\text{Tr}(R_{\text{ideal}}^{\dagger}R)+D}{D(D+1)}.$ (3)
Here, $D=2^{n}$ is the dimension of the computational Hilbert subspace ($n=2$), $R_{\text{ideal}}$ and $R$ are Pauli transfer matrices corresponding to the action of the ideal and simulated gates, respectively. Moreover, we calculate an average leakage out of the computational subspace $L=1-L_{e}$, where $L_{e}$ is a leakage error, and the mean projection of computational states population onto the target states after the CZ operation for each four states:
$\begin{split}P_{t}=\cfrac{1}{4}\sum_{\psi}|\langle\hat{\mathrm{CZ}}(\psi)|\hat{\mathcal{U}}(\psi)\rangle|^{2}=|\langle\psi|\hat{\mathcal{U}}(\psi)\rangle|^{2},\\ \psi\in\{|gg\rangle,|ge\rangle,|eg\rangle,|ee\rangle\}.\end{split}$ (4)
Here the unitary operator $\hat{\mathcal{U}}$ denotes the time evolution of the initial computational states during the gate implementation. The result is presented in Fig. 3(a), the corresponding gate duration as a function of an external flux amplitude is given in Fig. 3(b). Here we observe a periodic behavior of the gate fidelity. At the first maximum of the fidelity curve we do not get high-fidelity operation due to the immense ZZ interaction. Withal, the second maximum, marked with a star, shows the fidelity of 0.999(7) with 43.3-ns duration. For this point, a time evolution of the mean target state population (dashed blue), a conditional phase (red) and a leakage error (dashed-dotted green) are depicted in Fig. 3(c). The target state population behavior mostly relates to the oscillations between states during the gate implementation. The population leaks to the second excited qubit states $|gf\rangle,|fg\rangle$ and to the coupler state $|010\rangle$. The detailed dynamic of each computational state can be found in Appendix B.
Figure 4: The experimental benchmarking of the CZ gate. (a) The pulse sequence for the cross-entropy benchmarking (XEB) experiment with an interleaved CZ gate. (b) The depolarization fidelity averaged over 100 random circuits as a function of the circuit depth $m$ for the reference sequence: sequence (a) without CZ gates (red) and with an interleaved CZ operation (green). The blue dots corresponds to the square root of purity obtained by processing the interleaved XEB data with the speckle purity benchmarking technique. The dashed line is an exponential fit to the square root of purity. (c) The CZ gate quantum process tomography. The reconstructed Pauli transfer matrix gives the fidelity of $83.3\%$.
III.3 CZ gate implementation
We were able to experimentally implement the proposed CZ gate of the total 60-ns pulse duration with $-0.28$ V amplitude. The pulse edges length is $9$ ns, that is also included in the total gate duration. Since the experimental ZZ interaction strength is larger than the approximated one, the corresponding gate evolution in the experiment is longer. Thereby, we consider that the performed gate corresponds to the point nearby the third peak in Fig. 3(a). In order to validate the two-qubit gate performance we start with the quantum process tomography experiment, the result of which is depicted in Fig. 4(c). The fidelity of the reconstructed CZ Pauli transfer matrix is $83.3\%$. We explain the reduction of the gate quality by the poor multi-qubit readout. Therefore, we utilize cross-entropy benchmarking (XEB) [1] resilient against SPAM errors in order to characterize the performance of the implemented CZ operation.
For the XEB we execute the gate sequence shown in Fig. 4(a), where $C_{ij}$ are randomly chosen single-qubit Clifford gates. Here the first index corresponds to the qubit number and the second index refers to the serial number of the gate in the sequence. The single-qubit Clifford group is constructed of the $\pi/2$ rotations around X axis of $46.66$ ns duration and virtual S gates in the same way as it was performed in Ref. [44]. First, we accomplish reference XEB consisting only single-qubit operations simultaneously implemented on both qubits. The obtained depolarization fidelity averaged over randomly sampled quantum circuits as a function of a circuit depth $m$ is shown in Fig. 4(b) with a blue color. We approximate the data with the function $ap^{m}$, where $p$ is the depolarizing parameter and $a$ is a fitting parameter used to take into account a state preparation and measurement (SPAM) errors. The resulted depolarization parameter $p_{1}$ is $0.9714\pm 0.0005$. The similar data collected for the same circuits with the interleaved CZ gate is depicted in Fig. 4(b) in green. The depolarization parameter $p_{2}$ is equal to $0.947\pm 0.001$. In order to calculate the conventional fidelity $F$ of a target two-qubit gate we use the formula [1, 45]
$F=p+(1-p)/D,$ (5)
where $p=p_{2}/p_{1}$. The obtained CZ gate fidelity $F_{\mathrm{CZ}}$ is $(98.14\pm 0.12)\%$.
Since we use the XEB method, we can utilize it to estimate the acquired two-qubit phase $\theta$. By optimizing the linear cross-entropy between the experimental data and the simulated quantum circuits with an interleaved controlled-phase gate, we obtain $\theta=(1.03\pm 0.03)\pi$. Next, we leverage a speckle purity benchmarking to estimate the decoherence error [1]. The calculated square root of purity is $0.955\pm 0.001$. By comparing this value to $p_{2}$, it can be concluded that the major cause of the gate infidelity is a decoherence. Moreover, the CZ gate fidelity is higher than the single-qubit ones since CZ gate is shorter, see Table 1.
$f_{q}$ $T_{2}^{*}\;(\mu s)$ $T_{1}\;(\mu s)$ $t_{1Q}$ (ns)
Qubit A 6.450 8.6 15.8 $2\times 46.66$
Qubit B 6.494 10.8 8.3 $2\times 46.66$
Table 1: The device parameters: qubits frequency, dephasing and coherence times, and the length of single-qubit Clifford gates in Fig. 4(a). The coefficient 2 denotes the fact, that we construct the Clifford group out of two $R_{\mathrm{X}}(\pi/2)$ gates with duration of 46.66-ns.
IV Conclusion
[…]
Acknowledgments
The authors acknowledge Alexey Ustinov for fruitful discussions and useful comments on the manuscript. The work was supported by Rosatom in the framework of the Roadmap for Quantum computing (Contract No. 151/21-503, December 21, 2021 dated 12/21/2021).
Appendix A FULL-CIRCUIT HAMILTONIAN
The corresponding electrical circuit of the proposed two-qubit design with the three-mode coupler is shown in Fig. 5. We identified ten Josephson junctions $J_{i}$, where $i=1,\dots,10$, shunted by capacitances $C_{i}$ with $i=2,3,4,5,6$. The Josephson junction capacitances are denoted as $C_{Ji}$, where $i=1,2,3,8,9,10$. The coupler Josephson junction capacitances $C_{J4}$, $C_{J5}$, $C_{J6}$ and $C_{J7}$ are taken into account in corresponding total capacitances $C_{34}$, $C_{4}$, $C_{45}$. The mutual capacitances between the circuit nodes are indicated as $C_{ij}$, where $i$ and $j$ are the node numbers. The external fluxes can be applied to each SQUID of both qubits and the coupler.
The circuit parameter values used for the design are given in Table 2. This choice of parameters allows reaching the optimal range of an effective qubit-qubit interaction strength to have the interaction, tending to zero, and enable it to implement a fast two-qubit operation. All junctions are supposed to be fabricated by the shadow evaporation technique with critical current density $j=1~{}\mu\textnormal{A}/\mu\textnormal{m}^{2}$.
Figure 5: Equivalent lumped-element circuit for the proposed two-qubit system with a tunable coupler. Transmons (red) are connected through mutual capacitance with each other and with the coupler (blue). $C_{i}$ stands for the node capacitances with respect to the ground node, $C_{J_{i}}$ ($i=1,\dots,10$) are the capacitances of Josephson junctions, $C_{J1,2}=C_{J1}+C_{J2}$ $C_{J19,10}=C_{J9}+C_{J10}$ are total capacitances of SQUID junctions of two transmons, $C_{ij}$ are the mutual capacitances between nodes $i$ and $j$ that facilitate coupling between qubits. $J_{i}$ stands for Josephson inductance of the $i$-th junction. In order not to pile up the scheme, the coupler Josephson junction capacitances are not shown separately, but included in the capacitances between the corresponding nodes.
Josephson energy (GHz) Capacitance (fF) Capacitance (fF)
$E_{J1},\;E_{J10}$ 124.2 $C_{2}$ 75.6 $C_{35}$ 0.005
$E_{J2},\;E_{J9}$ 43.5 $C_{3}$ 47.2 $C_{56}$ 9.5
$E_{J3},\;E_{J8}$ 17.4 $C_{4}$ 144 $C_{J1},\;C_{J10}$ 11.1
$E_{J4},\;E_{J7}$ 124.2 $C_{5}$ 46.5 $C_{J2},\;C_{J9}$ 3.9
$E_{J5}$ 79.5 $C_{6}$ 75.9 $C_{J3},\;C_{J8}$ 1.5
$E_{J6}$ 129.1 $C_{23}$ 9 $C_{J4},\;C_{J7}$ 11.1
- - $C_{26}$ 0.015 $C_{J5}$ 11.5
- - $C_{34},\;C_{45}$ 13 $C_{J6}$ 7.1
Table 2: Target circuit parameters of the device: Josephson energies of junctions, node and Josephson junction capacitances. The coupler Josephson junction capacitances $C_{J4}$, $C_{J5}$ and $C_{J6}$, $C_{J7}$ are taken into account in corresponding values $C_{34}$, $C_{4}$, $C_{45}$.
Below, we describe each element of the device and derive the full system Hamiltonian.
The single-qubit Hamiltonian has the following form [38]
$\begin{split}\hat{H}_{\text{q}}&=4E_{\textnormal{C}_{q}}\hat{n}^{2}+E_{\textnormal{J3}}\left(1-\cos\hat{\varphi}_{2}\right)+\\ &+(E_{\textnormal{J1}}+E_{\textnormal{J2}})\cos\frac{\pi\Phi}{\Phi_{0}}\sqrt{1+d^{2}\operatorname{tan}^{2}\frac{\pi\Phi}{\Phi_{0}}}\left(1-\cos\hat{\varphi}_{1}\right),\end{split}$ (6)
where $\hat{\varphi}_{1}$, $\hat{\varphi}_{2}$ are the phase operators on junctions $J_{1}$, $J_{3}$ correspondingly, $\hat{n}$ is the Cooper pair number operator canonically conjugated to the sum $\hat{\varphi}=\hat{\varphi}_{1}+\hat{\varphi}_{2}$. The charge energy is expressed as $E_{\textnormal{C}_{q}}=e^{2}/2C_{\Sigma}$, where $C_{\Sigma}=C_{2}+C_{J3}C_{J1,2}/(C_{J3}+C_{J1,2})$. $E_{Ji}$ is Josephson energy of the $i$-th junction, $d=(E_{\textnormal{J2}}-E_{\textnormal{J1}})/(E_{\textnormal{J2}}+E_{\textnormal{J1}})$ is the SQUID asymmetry.
The coupler has three nodes, and therefore we obtain three coupler modes. Node phases can be conveniently expressed with the normal mode coordinates $\widetilde{\varphi}_{3},\widetilde{\varphi}_{4},\widetilde{\varphi}_{5}$:
$\varphi_{3}=\widetilde{\varphi}_{3}+\widetilde{\varphi}_{4};\quad\varphi_{4}=\widetilde{\varphi}_{3};\quad\varphi_{5}=\widetilde{\varphi}_{3}-\widetilde{\varphi}_{5}.$ (7)
Here and further, we omit the hats of quantum operators in order to avoid an overloading of the formulas. Hence, we have the coupler Hamiltonian
$\begin{split}H_{\text{c}}&=4E_{C_{00}}\widetilde{n}_{3}^{2}+4E_{C_{11}}\widetilde{n}_{4}^{2}+4E_{C_{22}}\widetilde{n}_{5}^{2}+2E_{C_{01}}\widetilde{n}_{3}\widetilde{n}_{4}+2E_{C_{02}}\widetilde{n}_{3}\widetilde{n}_{5}+2E_{C_{12}}\widetilde{n}_{4}\widetilde{n}_{5}+\\ &+E_{J4}\left(1-\cos\widetilde{\varphi}_{4}\right)+E_{J_{7}}\left(1-\cos\widetilde{\varphi}_{5}\right)+(E_{\textnormal{J5}}+E_{\textnormal{J6}})\cos\frac{\pi\Phi}{\Phi_{0}}\sqrt{1+d^{2}\textnormal{tan}^{2}\frac{\pi\Phi}{\Phi_{0}}}(1-\cos\widetilde{\varphi}_{3}).\end{split}$ (8)
Here $E_{C_{ij}}=e^{2}/2C_{ij}$, where $C_{ij}$ is the element of a $3\times 3$ capacitance matrix:
$\frac{1}{2}\begin{bmatrix}C_{\textnormal{3}}+C_{\textnormal{4}}+C_{\textnormal{5}}&C_{\textnormal{3}}&-C_{\textnormal{5}}\\ C_{\textnormal{3}}&C_{\textnormal{3}}+C_{\textnormal{34}}+C_{\textnormal{35}}&C_{\textnormal{35}}\\ -C_{\textnormal{5}}&C_{\textnormal{35}}&C_{\textnormal{35}}+C_{\textnormal{45}}+C_{\textnormal{5}}\end{bmatrix}.$ (9)
Eventually, the coupler Hamiltonian can be represented in the shortened form as three subsystems with capacitive interaction between each other:
$H_{\text{c}}=\sum_{i}H_{i}+\sum_{i\not=j}H_{ij},\quad i,j\in\{3,4,5\}.$ (10)
Here, the index $i$ corresponds to the Hamiltonian terms associated with the $i$-th phase coordinate.
We show the flux dependence of the coupler energy level structure obtained by numerical calculation in Fig. 6. There is an alternation of tunable and non-tunable energy levels that is due to the different contribution of the center node with the coupler SQUID.
For the proposed system we choose node fluxes $\varphi_{i}$, corresponding to nodes $i$ in Fig. 5, as the generalized coordinates of the system. We can write down the circuit Lagrangian $L(\varphi_{i},\dot{\varphi_{i}})$ using node fluxes together with the voltages $\dot{\varphi}_{i}$:
$L=T-U,$ (11)
$\begin{split}T=\frac{\Phi^{2}_{0}}{2(2\pi)^{2}}\Big{[}&C_{J1,2}\dot{\varphi}_{1}^{2}+C_{2}\dot{\varphi}_{2}^{2}+C_{\textnormal{J3}}(\dot{\varphi}_{2}-\dot{\varphi}_{1})^{2}+C_{\textnormal{23}}(\dot{\varphi}_{3}-\dot{\varphi}_{2})^{2}+C_{3}\dot{\varphi}_{3}^{2}+C_{\textnormal{34}}(\dot{\varphi}_{4}-\dot{\varphi}_{3})^{2}+\\ &+C_{4}\dot{\varphi}_{4}^{2}+C_{5}\dot{\varphi}_{5}^{2}+C_{\textnormal{45}}(\dot{\varphi}_{5}-\dot{\varphi}_{4})^{2}+C_{\textnormal{35}}(\dot{\varphi}_{5}-\dot{\varphi}_{3})^{2}+C_{6}\dot{\varphi}_{6}^{2}+\\ &+C_{\textnormal{56}}(\dot{\varphi}_{6}-\dot{\varphi}_{5})^{2}+C_{\textnormal{26}}(\dot{\varphi}_{6}-\dot{\varphi}_{2})^{2}+C_{\textnormal{J8}}(\dot{\varphi}_{7}-\dot{\varphi}_{6})^{2}+C_{J9,10}\dot{\varphi}_{7}^{2}\Big{]},\end{split}$ (12)
$\begin{split}U&=E_{\textnormal{J1}}\big{(}1-\cos\varphi_{1}\big{)}+E_{\textnormal{J2}}\big{(}1-\cos(\varphi_{1}-\varphi_{\text{ext},1})\big{)}+E_{\textnormal{J3}}\big{(}1-\cos(\varphi_{2}-\varphi_{1})\big{)}+\\ &+E_{\textnormal{J4}}\big{(}1-\cos(\varphi_{4}-\varphi_{3})\big{)}+E_{\textnormal{J5}}\big{(}1-\cos\varphi_{4}\big{)}+E_{\textnormal{J6}}\big{(}1-\cos(\varphi_{4}-\varphi_{\text{ext},c})\big{)}+\\ &+E_{\textnormal{J7}}\big{(}1-\cos(\varphi_{5}-\varphi_{4})\big{)}+E_{\textnormal{J8}}\big{(}1-\cos(\varphi_{6}-\varphi_{7})\big{)}+E_{\textnormal{J9}}\big{(}1-\cos\varphi_{7}\big{)}+\\ &+E_{\textnormal{J10}}\big{(}1-\cos(\varphi_{7}-\varphi_{\text{ext},2})\big{)},\end{split}$ (13)
where $T$ and $U$ are, respectively, the kinetic and potential energy. The kinetic energy term can be rewritten in matrix form $T=\frac{1}{2}\vec{\dot{\varphi}}^{T}C\vec{\dot{\varphi}}$, where $\vec{\dot{\varphi}}=[\dot{\varphi}_{1},\dot{\varphi}_{2},\dot{\varphi}_{3},\dot{\varphi}_{4},\dot{\varphi}_{5},\dot{\varphi}_{6},\dot{\varphi}_{7}]$ and $C$ is a $7\times 7$ capacitance matrix:
$C=\begin{bmatrix}C_{\textnormal{1$\Sigma$}}&-C_{\textnormal{J3}}&0&0&0&0&0\\ -C_{\textnormal{J3}}&C_{\textnormal{2$\Sigma$}}&-C_{23}&0&0&-C_{26}&0\\ 0&-C_{23}&C_{\textnormal{3$\Sigma$}}&-C_{\textnormal{34}}&-C_{35}&0&0\\ 0&0&-C_{\textnormal{34}}&C_{\textnormal{4$\Sigma$}}&-C_{45}&0&0\\ 0&0&-C_{35}&-C_{45}&C_{\textnormal{5$\Sigma$}}&-C_{\textnormal{56}}&0\\ 0&-C_{26}&0&0&-C_{56}&C_{\textnormal{6$\Sigma$}}&-C_{\textnormal{J8}}\\ 0&0&0&0&0&-C_{\textnormal{J8}}&C_{\textnormal{7$\Sigma$}}\\ \end{bmatrix},\$ (14)
where
$\begin{split}C_{\textnormal{1$\Sigma$}}&=C_{\textnormal{J1,2}}+C_{\textnormal{J3}},\\ C_{\textnormal{2$\Sigma$}}&=C_{\textnormal{2}}+C_{\textnormal{J3}}+C_{\textnormal{23}}+C_{\textnormal{26}},\\ C_{\textnormal{3$\Sigma$}}&=C_{\textnormal{3}}+C_{\textnormal{23}}+C_{\textnormal{34}}+C_{\textnormal{35}},\\ C_{\textnormal{4$\Sigma$}}&=C_{\textnormal{4}}+C_{\textnormal{34}}+C_{\textnormal{45}},\\ C_{\textnormal{5$\Sigma$}}&=C_{\textnormal{5}}+C_{\textnormal{45}}+C_{\textnormal{35}}+C_{\textnormal{56}},\\ C_{\textnormal{6$\Sigma$}}&=C_{\textnormal{6}}+C_{\textnormal{56}}+C_{\textnormal{26}}+C_{\textnormal{J8}},\\ C_{\textnormal{7$\Sigma$}}&=C_{\textnormal{J8}}+C_{\textnormal{J9,10}}.\end{split}$ (15)
Applying the coupler coordinate transformation (Eq. 7) to the capacitance matrix yields
$C_{\textnormal{new}}=T_{r}^{T}\times C\times T_{r},\quad$ (16)
where the transformation matrix $T_{r}$ is defined as:
$T_{r}=\frac{1}{2}\begin{bmatrix}1&0&0&0&0&0&0\\ 0&1&0&0&0&0&0\\ 0&0&1&1&0&0&0\\ 0&0&1&0&0&0&0\\ 0&0&1&0&-1&0&0\\ 0&0&0&0&0&1&0\\ 0&0&0&0&0&0&1\\ \end{bmatrix}.\$ (17)
The potential energy becomes
$\begin{split}U&=E_{\textnormal{J1}}\big{(}1-\cos\varphi_{1}\big{)}+E_{\textnormal{J2}}\big{(}1-\cos(\varphi_{1}-\varphi_{\text{ext},1})\big{)}+E_{\textnormal{J3}}\big{(}1-\cos(\varphi_{2}-\varphi_{1})\big{)}+\\ &+E_{\textnormal{J4}}\big{(}1-\cos\widetilde{\varphi}_{4}\big{)}+E_{\textnormal{J5}}\big{(}1-\cos\widetilde{\varphi}_{3}\big{)}+E_{\textnormal{J6}}\big{(}1-\cos(\widetilde{\varphi}_{3}-\varphi_{\text{ext},c})\big{)}+E_{\textnormal{J7}}\big{(}1-\cos\widetilde{\varphi}_{5}\big{)}+\\ &+E_{\textnormal{J8}}\big{(}1-\cos(\varphi_{6}-\varphi_{7})\big{)}+E_{\textnormal{J9}}\big{(}1-\cos\varphi_{7}\big{)}+E_{\textnormal{J10}}\big{(}1-\cos(\varphi_{7}-\varphi_{\text{ext},2})\big{)}.\end{split}$ (18)
Figure 6: Coupler energy levels as a function of the external magnetic flux in the coupler SQUID and corresponding system states. The coupler state is determined by the states of three subsystems. The three numbers reflect the number of excitations in each of the coupler modes. Levels are split due to the capacitive interaction between the nodes.
Figure 7: (a) Energy levels of the transmon-coupler-transmon system from external magnetic flux in the coupler with parameters from Tab.II. The system is determined by the states of five subsystems. The first symbol in the system quantum state corresponds to the state of the first transmon ($|g\rangle$, $|e\rangle$ or $|f\rangle$), the second symbol corresponds to the state of the second transmon (similarly to the first transmon), the rest corresponds to the state of the coupler as in Fig. 6. Coupler tunable mode pushes transmon levels, changing the value of ZZ-interaction between qubits. The system states that are important when considering population dynamics (see Fig. 8) are noted. (b) ZZ-interaction between qubits vs external magnetic flux on the coupler. The red circle shows the minimal value of ZZ-interaction.
By means of this transformation the coupler circuit can be divided into three subsystems. Thus, the matrices for calculation have a smaller dimension and calculations are faster. Taking into account two transmons, we have five subsystems with five levels in each one. We define grid for each phase $\varphi_{i}$ variable with phase step $\delta\varphi_{i}=\pi/16$ and corresponding canonically conjugate momenta $q_{i}=\frac{\partial L}{\partial{\dot{\varphi}}_{i}}.$ The grid for coupler external flux has the phase step $\delta\varphi_{ext}=\pi/128.$ The system Hamiltonian in terms of the first-order normal modes is defined as
$H=\sum_{i}q_{i}{\dot{\varphi}}_{i}-L=\frac{1}{2}\vec{q}^{T}C_{\text{new}}^{-1}\vec{q}+U,$ (19)
where $C_{\text{new}}^{-1}$ is the inverse capacitance matrix. We calculate the eigenvalues of each subsystem in phase basis using numerical diagonalization and find charge matrix elements by means of Fourier transformation. Then we construct the full Hamiltonian introducing the interaction terms in the form of Eq. (2) from the main text. Charge energy here is defined as $E_{Cij}=\frac{e^{2}}{2}C_{\text{new}}^{-1}.$ Finally, we numerically diagonalized the full Hamiltonian.
The calculated energy spectrum as a function of magnetic flux $\Phi$ in the coupler SQUID is plotted in Fig. 7(a). We provide the full system spectrum calculated for the qubits at the operating point as a function of an external magnetic flux $\Phi_{\text{ext}}$. The eigenstates are labelled as $|n_{\text{q1}},n_{\text{q2}},n_{\text{c1}},n_{\text{c2}},n_{\text{c3}}\rangle$, where $n_{i}$ is the occupancy of the $i$-th mode. The four computational levels are $|gg000\rangle$, $|eg000\rangle$, $|ge000\rangle$, $|ee000\rangle$.
As our goal is to implement a CZ gate, we are interested in the controllable ZZ interaction between the computational qubits. Here, the effective ZZ interaction is determined according to the conventional expression $\zeta_{ZZ}=f_{11}-f_{10}-f_{01}$. We numerically calculate the strength of the interaction for the transmon qubits, placed at upper sweet spots, and depict the result in Fig. 7(b). It can be seen that varying the external magnetic flux on the coupling element, one can change the effective interaction from 146 kHz up to 44 MHz. Thereby, the ZZ coupling is tuned off at the zero flux bias point ($\Phi=0$) and activated at the flux degeneracy point ($\Phi=0.5\Phi_{0}$).
$J_{1}$ $J_{2}$ $J_{3}$ $J_{4},J_{7}$ $J_{5}$ $J_{6}$ $J_{8}$ $J_{9}$ $J_{10}$
$E_{J}$ (GHz) 104.4 23.1 28.8 159.8 191.5 117.0 29.3 23.5 106.1
$C$ (fF) 9.3 2.1 2.6 14.2 17.1 10.4 2.6 2.1 9.5
Table 3: Josephson junction parameters extracted from the experiment: Josephson energy and capacitance
Figure 8: Time evolution of populations for four initial computational states during the implemented gate. The state notation corresponds to the mode occupations of the Hamiltonian. The flux pulse on the coupler has the shape shown in Fig. 3(d) in the main text. In this case, the fidelity of the CZ gate $F=0.999(7)$.
Appendix B TIME EVOLUTION OF STATE POPULATION
We investigated the time evolution of the computational states population during the CZ gate (see Fig. 8, corresponds to the star in Fig. 3(a) in the main text) with experimental device parameters (see Tab. 3). Besides exchange between four target states, we have leakages to the levels $|fg000\rangle$, $|gf000\rangle$ and $|gg010\rangle$. At the end of the two-qubit operation, the overall population returns to the computational subspace.
Appendix C EXPERIMENTAL SETUP
The experimental setup is presented in Fig. 9. The device is placed in the dilution refrigerator (grey area). Transmons are coupled to a readout reflection line via a resonator and a Purcell filter. Transmon and coupler control lines are attenuated and filtered for thermalization and noise reduction, as well as the readout line. We use an impedance-matching parametric amplifier [46] for the gain of a readout signal. Outside the refrigerator the system of microwave and dc control is assembled. For pulse generation we use IQ mixers with inputs for the sine signal generator and the arbitrary waveform generator. For the readout we use IQ mixer in the opposite direction: the amplified signal from the readout line comes to the RF port and converted to the I and Q signals, that go to the digitizer.
Figure 9: The scheme of the experimental setup.